Radio wave travel at the speed of light. Radio station KSTX broadcast at a frequency of 89.1 MHz. What is the wavelength of the radio station ratio wave


1
Expert's answer
2021-12-07T09:52:37-0500

Solution;

"\lambda =\frac{V}{f}=\frac{3×10^8}{89.1×10^6}=3.367m"